OpenStudy (anonymous):

Factor the trinomial: 4x2 - 8x - 5 A. (2x + 1)(2x - 5) B. (4x + 1)(x - 5) C. (4x - 5)(x + 1) D. prime

OpenStudy (amistre64):

you can always check by expanding the answers to see what matches

OpenStudy (amistre64):

it looks like the key element here is in the OI part of the fOIl method

OpenStudy (hoblos):

4x2 - 8x - 5 = 4x2 + 2x -10x -5 = 2x(2x+1) - 5(2x+1) = (2x+1)(2x-5)

OpenStudy (anonymous):